Prove that: sin4A-cos4A=sin2A-cos2A
1
Expert's answer
2012-09-14T12:09:41-0400
This statement is wrong. Let's take A = π/12, for example:

sin4A-cos4A = sin4(π/12)-cos4(π/12) = sin(π/3) - cos(π/3) = √3/2 - 1/2.

But

sin2A-cos2A = sin2(π/12)-cos2(π/12) = sin(π/6)-cos(π/6) = 0.5 - √3/2

and

& radic;3/2 - 1/2 ≠ 0.5 - √3/2.